# A Holistic Optimization Framework for Energy Efficient UAV-assisted Fog Computing: Attitude Control, Trajectory Planning and Task Assignment

### v2

@@ III-E3 Energy consumption for computation
$e_{ij}^{\mathrm{COMP}}(t)=o_{ij}^{\text{UAV}}(t)\delta_{u}s_{ij}(t)c_{ij}\left(f(t)\right)^{2}+o_{ij}^{\text{MD}}(t)\delta_{j}s_{ij}(t)c_{ij}\left(f_{j}(t)\right)^{2},$ (12)
where $\delta_{u}$ and $\delta_{j}$ are the effective computing capacities of the UAV, and the $j$th MD, respectively.
III-E4 Overall Energy Consumption
@@
where $N_{j}$ is the total number of tasks for the $j$th UD.
III-F Delay
@@ III-F1 Transmission delay
$d_{ij}^{\text{TR}}(t)$ $=o_{ij}^{\text{DC}}(t)\left(\frac{s_{ij}(t)}{C_{j}(t)r_{(u,D)}(t)}+\frac{s_{ij}(t)}{C_{j}(t)r_{(D,u)}(t)}\right)$ (14)
$+\left(o_{ij}^{\text{UAV}}(t)+o_{ij}^{\text{DC}}(t)\right)\left(\frac{s_{ij}(t)}{C_{j}(t)r_{(j,u)}(t)}+\frac{s_{ij}(t)}{C_{j}(t)r_{(u,j)}(t)}\right).$
III-F2 Computation delay
@@
$d_{ij}^{\mathrm{COMP}}(t)=o_{ij}^{\text{UAV}}(t)\frac{s_{ij}(t)c_{ij}}{f(t)}+o_{ij}^{\text{MD}}(t)\frac{s_{ij}(t)c_{ij}}{f_{j}(t)}.$ (15)
@@
III-F3 Queuing delay
Consider that the arrival rate for transmission from the $j$th MD to UAV is $\lambda_{j}$, and the processing rate at the UAV for receiving transmissions is $[L\cdot C_{j}(t)\cdot r_{j,u}(t)]/s_{j}(t)$, where $s_{j}(t)=\sum_{i=1}^{N_{j}}s_{ij}(t)$ is the sum of tasks data size from that MD. If the total arrival rate in larger than the processing rate, a queuing delay for transmission will be incurred [14], that is,
$d_{ij}^{\mathrm{Q}}(t)=\frac{\lambda_{j}\left(1-p_{j}\right)s_{j}(t)^{2}}{L\cdot C_{j}(t)r_{j,u}(t)\cdot\left(L\cdot C_{j}(t)r_{j,u}(t)-\lambda_{j}\left(1-p_{j}\right)s_{j}(t)\right)}.$ (16)
@@ III-F3 Queuing delay
$d_{ij}^{\mathrm{UAVQ}}(t)=\frac{\bar{Q}}{\sum_{j=1}^{K}\lambda_{j}\left(1-p_{j}\right)}-\frac{\sum_{j=1}^{K}\lambda_{j}\left(1-p_{j}\right)s_{j}(t)c_{j}}{\tau f(t)\sum_{j=1}^{K}\lambda_{j}\left(1-p_{j}\right)},$ (17)
where $\bar{Q}$ is a parameter related to the computing capacity of the UAV.
III-F4 Total delay
@@
IV Joint Optimization Problem
IV-A Attitude Control
IV-A1 Classical PID Control
@@
  • Integral control (I): an integral term is introduced for correction of the proportional control to maintain the system stability. Its coefficient will approach 00 as time goes by;
@@
[…]
where $k_{p}$, $k_{i}$ and $k_{d}$ are the proportional, integral and differential coefficients, respectively. The output of the PID control system, $U(t)$, will affect $\omega,\theta,\phi,\psi$, which will in turns have an impact on the power consumption of the UAV (see Section III-B and Appendix -A for details).
IV-A2 Fuzzy PID Control
A major shortcoming of the classical PID control is that, as the coefficients $k_{p}$, $k_{i}$, and $k_{d}$ are constants, it cannot adapt to dynamic environments and inter-dependent parameters as in UAV assisted fog computing scenarios. Therefore, we propose the fuzzy PID control algorithm, a nonlinear control technique with adaptive capability based on the Fuzzy set theory [20].
The fuzzy control process is demonstrated in Fig. 5. The key part of the process is the fuzzy controller at the top of Fig. 5), which determines $\Delta k_{p}$, $\Delta k_{i}$, and $\Delta k_{d}$, the adjustments in control coefficients, based on the yaw angle error $E$ and the error rate $EC$ from the previous estimation. The specific rules are shown in Table III. Based on the demand for Mamdani reasoning method in solving fuzzy PID problems [46], we establish the the allowable value of $E$, $EC$, $\Delta k_{p}$, $\Delta k_{i}$, and $\Delta k_{d}$ in Table IV.
TABLE III: FUZZY PID CONTROL RULES
NB: Negative Big ($-3$)  NM: Negative Middle ($-2$)   NS: Negative Small ($-1$)  ZO (0): zero
@@ IV-A2 Fuzzy PID Control
[…]
@@
TABLE IV: FUZZY PID PARAMETER SETTINGS
@@
$E$, $EC$6, -5, -4, -3, -2, -3, -2, -1, -3, -2, -1,
-1, 0, 1, 2, $\Delta k_{p}$, $\Delta k_{d}$ 0, 1, $\Delta k_{i}$ 0, 1,
3, 4, 5, 6 2, 3 2, 3
$\left\{\begin{array}[]{l}k_{p-Fuzzy}=k_{p}+\Delta k_{p}\\ k_{i-Fuzzy}=k_{i}+\Delta k_{i}\\ k_{d-Fuzzy}=k_{d}+\Delta k_{d}.\end{array}\right.$ (20)
Figure 5: Schematic structure of Fuzzy PID control system for quadrotor UAV.
IV-B Joint optimization model
$\begin{array}[]{ll}\underset{\{\mathbf{o}_{ij}(t),\mathbf{v}(t),\omega_{i}(t)|t\in\{0,\cdots,T\}\}}{\operatorname{Min}}&\mathbb{S}=\sum\limits_{t=0}^{T}[\mathbb{D}(t)+\epsilon\cdot\mathbb{E}(t)]\\ \text{s.t.}&\left(\complement_{1}\right)\hskip 5.69046pt0\leq|\mathbf{v}(t)|\leq v^{\max}\\ &\left(\complement_{2}\right)\hskip 5.69046pt0\leq z(t)\leq z^{\max}\\ &\left(\complement_{3}\right)\hskip 5.69046pt\sum_{j=1}^{K}C_{j}(t)\leq N_{c}\\ &\left(\complement_{4}\right)\hskip 5.69046ptf^{\min}\leq f(t)\leq f^{\max}\\ &\left(\complement_{5}\right)\hskip 5.69046ptf_{j}^{\min}\leq f_{j}(t)\leq f_{j}^{\max}\\ &\left(\complement_{6}\right)\hskip 5.69046ptp^{\min}\leq p(t)\leq p^{\max}\\ &\left(\complement_{7}\right)\hskip 5.69046ptp_{j}^{\min}\leq p_{j}(t)\leq p_{j}^{\max}\\ &\left(\complement_{8}\right)\hskip 5.69046pto_{ij}^{\text{MD}}(t)+o_{ij}^{\text{UAV}}(t)+o_{ij}^{\text{DC}}(t)=1\\ &\left(\complement_{9}\right)\hskip 5.69046pto_{ij}^{\text{MD}}(t),o_{ij}^{\text{UAV}}(t),o_{ij}^{\text{DC}}(t)\in\{0,1\}\\ &\left(\complement_{10}\right)\sum_{t=0}^{T}\left(E^{\text{UAV}}(t)\right)\leq B_{c}\\ &i\in\{1,\ldots,N_{j}\},j\in\{1,\ldots,K\}\end{array}$ (21)
The decision variables include the angular velocities $\omega_{i}(t)$ related to attitude control, the velocity vector $\mathbf{v}(t)$ for trajectory planning, and task assignment decisions $\mathbf{o}_{ij}(t)$. The parameter $\epsilon$ is a weighting factor in the objective function, which flexibly caters to different magnitudes and ranges of the two measurements, as well as reflects the preference in terms of the trade-off between the two metrics in various fog computing scenarios. We will demonstrate in the result section that our proposed approach can achieve high performances for a wide range of $\epsilon$ values. In terms of the constraints, $\complement_{1}$ to $\complement_{3}$ were described in detail previously when they first appeared. $\complement_{4}$ to $\complement_{7}$ ensure that the processing frequency and transmission power of UAV and MDs are always within the effective range. $\complement_{8}$ and $\complement_{9}$ guarantee that every task is executed at exactly one place at any moment (no duplicate computing). $\complement_{10}$ restricts that the total energy consumption of the UAV during $T$ does not exceed its battery capacity, as $E^{\text{UAV}}(t)$ denotes the energy consumption of the UAV at the $t$th timeslot,
$E^{\text{UAV}}(t)$ $=E^{\text{MOV}}(t)+\sum_{j=1}^{K}\sum_{i=1}^{N_{j}}\left[o_{ij}^{\text{DC}}(t)\left(\frac{p_{u}(t)s_{ij}(t)}{C_{j}(t)r_{(u,D)}(t)}\right)\right.$ (22)
$\left.+o_{ij}^{\text{UAV}}(t)\left(\frac{p_{u}(t)s_{ij}(t)}{C_{j}(t)r_{(u,j)}(t)}+\delta_{u}s_{ij}(t)c_{ij}\left(f(t)\right)^{2}\right)\right].$
Note that the processing frequencies $f(t)$, $f_{j}(t)$, and transmission powers $p(t)$, $p_{j}(t)$ are dependent on the decision variables $\mathbf{o}_{ij}(t)$. Numerous scholarly works provide comprehensive analyses of these intricate interactions, e.g., [14].
V Algorithms
The problem (21) possesses an intrinsic nonconvex nature. We divide the problem into two subproblems, one focusing on generating an initial trajectory and altitude, and the other on task assignment and resource allocation, as well as further tuning the trajectory. We propose two algorithms to address the two subproblems, respectively, with enhanced decoupling and safety values mechanism incorporated in the second algorithm to overcome the deadlock issue in conventional heuristic algorithms.
An overview of the steps in our proposed framework to address the joint optimization problem is shown in Fig. 6. We will apply an Ant Colony System based algorithm (Algorithm 2, to be described in detail later) to determine an initial trajectory based on the terrain of the area and the position of MDs. Then, the attitude control mechanism described in Section IV-A will be invoked to determine the as well as tune the trajectory. Finally, a Particle Swarm Optimization based algorithm (Algorithm 1, which was initially proposed in the conference version [15]) will determine the optimal task assignments and resource allocations. We will also analyze the convergence performance of the proposed algorithms at the end of the section.

### v3

@@ 3.5.3 Energy consumption for computation
$e_{ij}^{\mathrm{COMP}}(t)=o_{ij}^{\text{UAV}}(t)\delta_{u}s_{ij}(t)c_{ij}\left(f(t)\right)^{2}+o_{ij}^{\text{MD}}(t)\delta_{j}s_{ij}(t)c_{ij}\left(f_{j}(t)\right)^{2},$
[…]
3.5.4 Overall Energy Consumption
@@
where $N_{j}$ is the total number of tasks the $j$th MD.
3.6 Delay Model
@@ 3.6.1 Transmission delay
$\begin{aligned} d_{ij}^{\text{TR}}(t)&=o_{ij}^{\text{DC}}(t)\left(\frac{s_{ij}(t)}{C_{j}(t)r_{(u,D)}(t)}+\frac{s_{ij}(t)}{C_{j}(t)r_{(D,u)}(t)}\right)\\ &\quad+\left(o_{ij}^{\text{UAV}}(t)+o_{ij}^{\text{DC}}(t)\right)\left(\frac{s_{ij}(t)}{C_{j}(t)r_{(j,u)}(t)}+\frac{s_{ij}(t)}{C_{j}(t)r_{(u,j)}(t)}\right).\end{aligned}$
3.6.2 Computation delay
@@
$d_{ij}^{\mathrm{COMP}}(t)=o_{ij}^{\text{UAV}}(t)\frac{s_{ij}(t)c_{ij}}{f(t)}+o_{ij}^{\text{MD}}(t)\frac{s_{ij}(t)c_{ij}}{f_{j}(t)}.$
where $f(t)$ (unit: GHz) denotes the UAV’s processing frequency, bounded by $f^{\min}=1.0$ GHz and $f^{\max}=3.0$ GHz, representing a high-performance onboard processor. Similarly, $f_{j}(t)$ (unit: GHz) denotes the $j$th MD’s processing frequency, bounded by $f_{j}^{\min}=0.5$ GHz and $f_{j}^{\max}=2.0$ GHz. The UAV distributes its processing frequency $f(t)$ among all assigned tasks according to the joint optimization algorithm, which minimizes the overall weighted energy–delay cost. Thus, resource allocation is implicitly determined by the solution of the optimization problem rather than by a fixed scheduling rule.
@@
3.6.3 Queuing delay
We consider two types of queuing delays: transmission queuing and computation queuing. Transmission queuing occurs when the arrival rate exceeds the communication service rate, while computation queuing occurs when the aggregated task workload exceeds the UAV’s effective computing rate.
Our queuing model adopts an M/M/1-type abstraction, which is a standard first-order modeling approach in MEC literature for capturing queuing behavior under stochastic task arrivals [51, 20]. This abstraction enables tractable analysis while providing reasonable approximations of system performance under moderate to high load conditions. While more detailed queueing network models (e.g., multi-server queues or priority-based scheduling) could offer finer granularity, they would significantly complicate the optimization problem formulation. Our simplified model remains compatible with such extensions, which could be incorporated as future work when detailed scheduling policies are required.
Consider that the arrival rate for transmission from the $j$th MD to UAV is $\lambda_{j}$, and the processing rate at the UAV for receiving transmissions is $L\cdot C_{j}(t)\cdot r_{j,u}(t)]/s_{j}(t)$, where $s_{j}(t)=\sum_{i=1}^{N_{j}}s_{ij}(t)$ is the sum of tasks data size from that MD. Let $\pi_{j}$ denote the probability that an arriving task from the $j$th MD does not enter the transmission/computation queue (e.g., executed locally), so the effective arrival rate is $\lambda_{j}(1-\pi_{j})$. If the total arrival rate is larger than the processing rate, a queuing delay for transmission will be incurred [39], that is,
$d_{ij}^{\mathrm{Q}}(t)=\frac{\lambda_{j}\left(1-\pi_{j}\right)s_{j}(t)^{2}}{L\cdot C_{j}(t)r_{j,u}(t)\cdot\left(L\cdot C_{j}(t)r_{j,u}(t)-\lambda_{j}\left(1-\pi_{j}\right)s_{j}(t)\right)}.$
@@ 3.6.3 Queuing delay
$d_{ij}^{\mathrm{UAVQ}}(t)=\frac{\bar{Q}}{\sum_{j=1}^{K}\lambda_{j}\left(1-\pi_{j}\right)}-\frac{\sum_{j=1}^{K}\lambda_{j}\left(1-\pi_{j}\right)s_{j}(t)c_{j}}{\tau f(t)\sum_{j=1}^{K}\lambda_{j}\left(1-\pi_{j}\right)},$
where $\bar{Q}=5\times 10^{7}$ (unit: bit-equivalent) is a queue-capacity parameter representing a queue threshold corresponding to the UAV’s onboard computing throughput, and $\tau=1$ s is a normalization constant that normalizes the aggregated computational workload into an effective service time per task.
3.6.4 Total delay
@@
Congestion in the network manifests through the growth of queue lengths, which directly contributes to the queuing delay components $d_{ij}^{\mathrm{Q}}(t)$ and $d_{ij}^{\mathrm{UAVQ}}(t)$ in the total delay expression. Since our objective function minimizes the weighted sum of delay and energy consumption, congested conditions (characterized by increased queuing delays) are implicitly penalized. This approach eliminates the need for an explicit congestion metric, as the optimization naturally steers the system away from congestion-prone configurations to minimize the overall operational cost.
4 Joint Optimization Problem
$\begin{array}[]{ll}\underset{\{\mathbf{o}_{ij}(t),\mathbf{v}(t),\omega_{i}(t)|t\in\{0,\cdots,T\}\}}{\operatorname{Min}}&\mathbb{S}=\sum\limits_{t=0}^{T}[\mathbb{D}(t)+\epsilon\cdot\mathbb{E}(t)]\\ \text{s.t.}&{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\left(\complement_{1}\right)\hskip 5.69046pt\omega^{\min}\leq\omega_{i}(t)\leq\omega^{\max},\quad\forall i}}\\ &\left(\complement_{2}\right)\hskip 5.69046pt0\leq|\mathbf{v}(t)|\leq v^{\max}\\ &\left(\complement_{3}\right)\hskip 5.69046pt0\leq z(t)\leq z^{\max}\\ &\left(\complement_{4}\right)\hskip 5.69046pt\sum_{j=1}^{K}C_{j}(t)\leq N_{c}\\ &\left(\complement_{5}\right)\hskip 5.69046ptf^{\min}\leq f(t)\leq f^{\max}\\ &\left(\complement_{6}\right)\hskip 5.69046ptf_{j}^{\min}\leq f_{j}(t)\leq f_{j}^{\max}\\ &\left(\complement_{7}\right)\hskip 5.69046ptp^{\min}\leq p_{u}(t)\leq p^{\max}\\ &\left(\complement_{8}\right)\hskip 5.69046ptp_{j}^{\min}\leq p_{j}(t)\leq p_{j}^{\max}\\ &\left(\complement_{9}\right)\hskip 5.69046pto_{ij}^{\text{MD}}(t)+o_{ij}^{\text{UAV}}(t)+o_{ij}^{\text{DC}}(t)=1\\ &\left(\complement_{10}\right)\hskip 5.69046pto_{ij}^{\text{MD}}(t),o_{ij}^{\text{UAV}}(t),o_{ij}^{\text{DC}}(t)\in\{0,1\}\\ &\left(\complement_{11}\right)\hskip 5.69046pt\sum_{t=0}^{T}\left(E^{\text{UAV}}(t)\right)\leq B_{c}\\ &i\in\{1,\ldots,N_{j}\},j\in\{1,\ldots,K\}\end{array}$
The decision variables include the angular velocities $\omega_{i}(t)$ related to attitude control, the velocity vector $\mathbf{v}(t)$ for trajectory planning, and task assignment decisions $\mathbf{o}_{ij}(t)$. The parameter $\epsilon$ is a weighting factor in the objective function, which flexibly caters to different magnitudes and ranges of the two measurements, as well as reflects the preference in terms of the trade-off between the two metrics in various fog computing scenarios. We will demonstrate in the result section that our proposed approach can achieve high performances for a wide range of $\epsilon$ values. In terms of the constraints, $\complement_{1}$ to $\complement_{4}$ are constraints imposed by the hardware limitations of the respective components. $\complement_{5}$ to $\complement_{8}$ ensure that the processing frequency and transmission power of UAV and MDs are always within the effective range. $\complement_{9}$ and $\complement_{10}$ guarantee that every task is executed at exactly one place at any moment (no duplicate computing). $\complement_{11}$ restricts that the total energy consumption of the UAV during $T$ does not exceed its battery capacity, as $E^{\text{UAV}}(t)$ denotes the energy consumption of the UAV at the $t$th timeslot,
$\begin{aligned} E^{\text{UAV}}(t)&=E^{\text{MOV}}(t)+\sum_{j=1}^{K}\sum_{i=1}^{N_{j}}\left[o_{ij}^{\text{DC}}(t)\left(\frac{p_{u}(t)s_{ij}(t)}{C_{j}(t)r_{(u,D)}(t)}\right)\right.\\ &\left.+o_{ij}^{\text{UAV}}(t)\left(\frac{p_{u}(t)s_{ij}(t)}{C_{j}(t)r_{(u,j)}(t)}+\delta_{u}s_{ij}(t)c_{ij}\left(f(t)\right)^{2}\right)\right].\end{aligned}$
Note that the processing frequencies $f(t)$, $f_{j}(t)$, and transmission powers $p_{u}(t)$, $p_{j}(t)$ are dependent on the decision variables $\mathbf{o}_{ij}(t)$. Numerous scholarly works (e.g., [39]) provided comprehensive analyses of these intricate interactions.
5 Algorithms
The problem (19) possesses an intrinsic nonconvex nature. We divide the problem into three subproblems, one focusing on controlling the attitude, the second on generating an initial trajectory and altitude, and the other on task assignment and resource allocation, as well as further tuning the trajectory, respectively.
5.1 Attitude Control
5.1.1 Classical PID Control
@@
  • 2.
    Integral control (I): an integral term is introduced for correction of the proportional control to maintain the system stability. Its coefficient will approach 0 as time goes by;
  • 3.
@@
$U(t)=k_{p}e(t)+k_{i}\int e(t)dt+k_{d}\frac{d}{e}(t){dt},$ (21)
where $k_{p}$, $k_{i}$ and $k_{d}$ are the proportional, integral and differential coefficients, respectively. The output of the PID control system, $U(t)$, will affect $\omega,\theta,\phi,\psi$, which will in turns have an impact on the power consumption of the UAV (see Section 3.2 and A for details).
5.1.2 Fuzzy PID Control
The fuzzy PID control algorithm improves on the classical PID control by adding nonlinear adaptive capability through Fuzzy Set Theory [46, 7]. By inputting fuzzification and a set of fuzzy rules, the PID gain is dynamically adjusted according to the current error $E$ and the error rate $EC$. In addition, the coefficients $k_{p}$, $k_{i}$, and $k_{d}$ are no longer constants, enabling them to adapt to environmental changes. However, since the fuzzy rules are preset based on expert knowledge, the controller lacks flexibility in adapting to unforeseen complex environmental changes in real time [2].
5.1.3 FEAR-PID Control
To solve the above problems, we propose the Fuzzy-Enhanced Adaptive Reinforcement PID (FEAR-PID) controller. FEAR-PID integrates RL in fuzzy PID to provide an additional adaptive layer. The control process of FEAR-PID is demonstrated in Fig. 5. Specifically, the initial control strategy is first provided by fuzzy logic based on expert knowledge. The specific rules are shown in Table 3.
Table 3: INITIAL FUZZY PID CONTROL RULES
NB: Negative Big (3-3)  NM: Negative Middle (2-2)   NS: Negative Small (1-1)  ZO (0): zero
@@ 5.1.3 FEAR-PID Control
[2pt/2pt] $\Delta k_{p}$$\Delta k_{i}$$\Delta k_{d}$ $\Delta k_{p}$$\Delta k_{i}$$\Delta k_{d}$ $\Delta k_{p}$$\Delta k_{i}$$\Delta k_{d}$ $\Delta k_{p}$$\Delta k_{i}$$\Delta k_{d}$ $\Delta k_{p}$$\Delta k_{i}$$\Delta k_{d}$ $\Delta k_{p}$$\Delta k_{i}$$\Delta k_{d}$ $\Delta k_{p}$$\Delta k_{i}$$\Delta k_{d}$
@@
Then, the PID gain coefficients in the fuzzy controller are obtained by a Double Deep Q-Network (DDQN). The RL module continuously optimizes the PID gain coefficients based on real-time feedback, thus effectively adapting to the dynamic environment and improving the control effect. The RL agent in FEAR-PID is designed as follows:
  • 1.
    State: The state $S_{t}$ at timeslot $t$ consists of the $E$, ECEC, and the PID coefficients $\Delta k_{p}$, $\Delta k_{i}$, and $\Delta k_{d}$. Represented as $S_{t}=[E,EC,\Delta k_{p},\Delta k_{i},\Delta k_{d}]$.
  • 2.
    Action: The action $A_{t}=[\Delta k_{p},\Delta k_{i},\Delta k_{d}]$ represents the adjustments made to the PID coefficients at each step, resulting in updated PID gains,
    $\left\{\begin{array}[]{l}k_{p-FEAR}=k_{p}+\Delta k_{p}\\ k_{i-FEAR}=k_{i}+\Delta k_{i}\\ k_{d-FEAR}=k_{d}+\Delta k_{d}.\end{array}\right$. (22)
  • 3.
    Reward: The reward function $R_{t}$ is designed to guide the controller toward minimizing the control error and achieving stable responses.
    $R_{t}=-\left(\vartheta_{1}|E|+\vartheta_{2}\left|\frac{dE}{dt}\right|+\vartheta_{3}\cdot\Delta U^{2}+\vartheta_{4}\cdot\Upsilon\right)$ (23)
    where the absolute error $|E|$ aims to minimize the error. The term $\left|\frac{dE}{dt}\right|$ controls the rate of error change for stability, $\Delta U^{2}$ captures control output variation, incentivizing smoother adjustments, AND $\Upsilon$ denotes the overshoot penalizing excessive responses and aids in maintaining stability. The weights $\vartheta_{1},\vartheta_{2},\vartheta_{3},$ and $\vartheta_{4}$ are provided to balance accuracy, convergence speed, stability, and control smoothness. The specific values and ranges of relevant parameters summarized in Table 4.
Table 4: FEAR-PID PARAMETER SETTINGS
@@
$E$, ECEC -6, -5, -4, -3, -2, -1, $\Delta k_{p}$, $\Delta k_{d}$ -3, -2, -1, 0, $\Delta k_{i}$ -1.5, -1, 0,
0, 1, 2, 3, 4, 5, 6 1, 2, 3 1, 1.5
$\vartheta_{1}$ 3 $\vartheta_{2}$ 0.5 $\vartheta_{3},\vartheta_{4}$ 1.5
Figure 5: Schematic structure of FEAR-PID control system for UAV.
5.2 Task Assignment and Resource Allocation
An overview of the steps in our proposed framework to address the task assignment, resource allocation and trajectory planning modules in the joint optimization problem is shown in Fig. 6. We will apply an Ant Colony System based algorithm (Algorithm 2, to be described in detail later) to determine an initial trajectory based on the terrain of the area and the position of MDs. Then, the attitude control mechanism described in Section 5.1 will be invoked to determine the as well as tune the trajectory. Finally, a Particle Swarm Optimization based algorithm (Algorithm 1, which was initially proposed in the conference version [24]) will determine the optimal task assignments and resource allocations. We will also analyze the convergence performance of the proposed algorithms at the end of the section.